what number is three more than twice the other number if the sum of the two numbers is 21
Answer by algebrahouse.com(1659)

"what number is three more than twice the other number if the sum of the two numbers is 21"

x = one number
2x + 3 = other number {one number is three more than twice the other}

x + 2x + 3 = 21 {their sum is 21}
3x + 3 = 21 {combined like terms}
3x = 18 {subtracted 3 from each side}
x = 6 {divided each side by 3}

2x + 3 {other number}
= 2(6) + 3 {substituted 6, in for x, into 2x + 3}
= 12 + 3 {multiplied}
= 15 {added}

one number is 6
other number is 15

For more help from me, visit: www.algebrahouse.com